# Minimal Construction of Graphs with Maximum Robustness

### v1

@@ 7.2 Consensus with an Edge Removal
These $\gamma$- and $(\gamma,\gamma)$ graphs achieve their respective maximum robustness levels using the fewest edges possible for a given number of nodes. In other words, removing any edge reduces their robustness levels at least by one. To validate this minimalistic property, we compare the consensus performance of the original $\gamma$- and $(\gamma,\gamma)$-MERGs with that of modified versions in which a single edge has been removed. For each graph, comparison is conducted using identical initial states for the normal agents and in the presence of $F$ Byzantine agents.
@@ 7.2.1 Simulations with $\gamma$-MERGs
We construct $5$-MERGs with $n=9$ and $n=10$ nodes as shown in Figure 2. Note these graphs achieve $5$-robustness with the fewest edges possible for a given number of nodes, and they can tolerate at most $F=2$ byzantine agents. We set agents $i\in\{0,1\}$ as Byzantine, and agents $i\in\{2,\dots,n-1\}$ as normal. Each normal agent $i$ is initialized with a state $x_{i}[0]\in\mathbb{R}$ sampled uniformly at random from the interval $[15,100]$ for $i\in\{2,\dots,5\}$, from $[0,7]$ for $i\in\{6,\dots,n-2\}$, and from $[8,14]$ for $i=n-1$. For all $t\geq 0$, all Byzantine agents adversarially share a value of $100$ with agents $i=0,\dots,\lceil\frac{n}{2}\rceil$ and of 00 with other remaining agents.
Figure 6 (a)-(b) and (e)-(f) illustrate the consensus results with and without edge removal. We have removed the edge $(3,8)$ for $n=9$ and $(4,9)$ for $n=10$ from the $5$-MERGs shown in Figure 2. In the $5$-MERGs, normal agents successfully reach resilient consensus (shown in Figure 6 (e) and (f)). However, removing the edge $(3,8)$ from each $5$-MERG reduces its robustness to $r=4$, which is insufficient to guarantee resilient consensus in the presence of $F=2$ Byzantine agents. As a result, normal agents in the modified graph fail to reach resilient consensus (as shown by the red arrows in Figure 6 (a) and (b)).
@@ 7.2.2 Simulations with $(\gamma,\gamma)$-MERGs
We use Theorem 6.26 to construct $(5,5)$-MERGs with $n=9$ and $n=10$ agents, as shown in Figure 4. These graphs can tolerate up to $F=4$ Byzantine agents. We set agents $i\in\{0,\dots,3\}$ as Byzantine, and agents $i\in\{4,\dots,n-1\}$ as normal. Each normal agent $i$ is initialized with a state $x_{i}[0]\in\mathbb{R}$ sampled uniformly at random from the interval $[50,100]$ for $i\in\{4,\dots,n-2\}$ and from $[1,50]$ for $i=n-1$. Byzantine agents share adversarial values: agent $3$ transmits $100$ to its neighbors at all times, while the remaining Byzantine agents transmit 00 to their neighbors $\forall t\geq 0$.
Figure 6 (c)-(d) and (g)-(h) illustrate the consensus results with and without edge removal. We have removed the edge $(7,8)$ for $n=9$ and $(0,2)$ for $n=10$ from the $(5,5)$-MERGs visualized in Figure 4. In these $(5,5)$-MERGs, normal agents successfully reach resilient consensus (shown in Figure 6 (g) and (h)). However, removing the edge from each $(5,5)$-MERG reduces its robustness to at most $(5,4)$-robustness, which is insufficient to guarantee resilient consensus in the presence of $F=4$ Byzantine agents. As a result, normal agents fail to reach resilient consensus in the modified graph (as shown by the red arrows in Figure 6 (c) and (d)).
@@ 9 Appendix
Here, the proof for Lemma 3.7 is provided below, accompanied by Figure 7 for illustrative support.
Figure 7: These figures demonstrate the first two iterations of the first part in the proof of Lemma 3.7. Throughout figures (a) to (d), we maintain $|\mathcal{S}_{1}|=\gamma-1$ and $|\mathcal{S}_{2}|=\gamma+1$. Figures (a) and (b) illustrate the first iteration. Here, $\mathcal{S}_{2}$ contains $\gamma+1$ nodes including $k$-clique $\mathcal{C}_{k,1}=\{v_{1},\cdots,v_{k}\}$ (colored in green), while $\mathcal{S}_{1}$ contains the remaining $\gamma-1$ nodes. Since $\mathcal{S}_{1}$ has to be $\gamma$-reachable, there exists a node $u_{1}\in\mathcal{S}_{1}$ that has edges with at least $\gamma$ nodes in $\mathcal{S}_{2}$. Figure (a) shows the case scenario where the yellow node $u_{1}$ does not have an edge with $v_{1}$ (highlighted in yellow). Then, $u_{1}$ forms a $k$-clique $\mathcal{C}_{k,2}$ with $\mathcal{C}_{k,1}\setminus\{v_{1}\}$. Consequently, in any future iteration, if any node $u_{i}\in\mathcal{S}_{1}$, $i\neq 1$, connects to at least $\gamma$ nodes in $\mathcal{S}_{2}$ but not to $v_{1}$, it will form a $(k+1)$-clique with $\mathcal{C}_{k,2}$. Then, $u_{1}$ gets swapped with a node in $S_{2}\setminus\mathcal{C}_{k,1}$, as shown in Figure (b). Figures (c) and (d) present the second iteration. Figure (c) presents the case scenario where $u_{2}$ only has $\gamma$ edges and does not have an edge with $v_{2}$ (highlighted in red). As before, if any node $u_{i}\in\mathcal{S}_{2}$, $i\neq 2$, connects to $\gamma$ nodes while missing an edge to $v_{2}$ in future iteration, it will form a $(k+1)$-clique. Then, $u_{2}$ gets swapped with a node in $\mathcal{S}_{2}\setminus U^{\prime}$ where $U^{\prime}=\mathcal{C}_{k,1}\cup\mathcal{C}_{k,2}$, as shown on Figure (d). The process continues, maintaining $\gamma$-reachability of $\mathcal{S}_{1}$, until a $\big{(}\lfloor\frac{\gamma+3}{2}\rfloor\big{)}$-clique is formed.
Proof 9.29.
@@
Second Part: Continuing from the previous paragraph, at $k=\lfloor\frac{\gamma+1}{2}\rfloor$ and for even values of $\gamma$, at worst the case 2) mentioned above can be repeated $\frac{\gamma}{2}$ times. Then, $\mathcal{S}_{2}$ contains $\mathcal{K}=\{u_{1},\cdots,u_{\frac{\gamma}{2}}\}$, $\mathcal{C}_{k,1}=\{v_{1},\cdots,v_{\frac{\gamma}{2}}$}, and a node $p$ not in $\mathcal{K}$ and $\mathcal{C}_{k,1}$. Note both $\mathcal{K}$ and $\mathcal{C}_{k,1}$ disjointly form a $\big{(}\frac{\gamma}{2}\big{)}$-clique. We know that a node $u_{m}\in\mathcal{S}_{1}$, $m=\frac{\gamma+2}{2}$, must form a $m$-clique $\mathcal{Q}$ with either $\mathcal{K}$ or $\mathcal{C}_{k,1}$. We swap $u_{m}\in\mathcal{S}_{1}$ with $p\in\mathcal{S}_{2}$. Now let $\mathcal{P}=\mathcal{S}_{2}\setminus\mathcal{Q}$. Note that nodes in $\mathcal{P}$ form a $(\frac{\gamma}{2})$-clique such that $\mathcal{P}\cap\mathcal{Q}=\emptyset$. Since $\mathcal{S}_{1}$ is $\gamma$-reachable even after $u_{m}$ is swapped out, $\exists u_{m+1}\in\mathcal{S}_{1}$ such that $|\mathcal{N}_{u_{m+1}}\cap\mathcal{S}_{2}|\geq\gamma$. There are two cases:
@@
If 1), $\mathcal{Q}\cup\{u_{m+1}\}$ forms a $\big{(}\frac{\gamma+4}{2}\big{)}$-clique. If 2), $\mathcal{P}\cup\{u_{m+1}\}$ forms a $\big{(}\frac{\gamma+2}{2}\big{)}$-clique $\mathcal{P}^{\prime}$ such that $\mathcal{P}^{\prime}\cap\mathcal{Q}=\emptyset$. In this case, WLOG, let $\mathcal{P}^{\prime}\subset\mathcal{S}_{1}$ and $\mathcal{Q}\subset\mathcal{S}_{2}$ such that $|\mathcal{S}_{1}|=|\mathcal{S}_{2}|=\gamma$. Since either $\mathcal{S}_{1}$ or $\mathcal{S}_{2}$ is $\gamma$-reachable, all of the nodes in either $\mathcal{P}^{\prime}$ or $\mathcal{Q}$ have edges with one additional node in $\mathcal{S}_{2}$ or $\mathcal{S}_{1}$ respectively, forming a $\big{(}\frac{\gamma+4}{2}\big{)}$-clique. Thus, $\mathcal{G}$ must contain a $\big{(}\frac{\gamma+4}{2}\big{)}$-clique for even $\gamma$. Since $\mathcal{G}$ must have a clique size of $\frac{\gamma+3}{2}$ and $\frac{\gamma+4}{2}$ for odd and even $\gamma$ respectively, $\mathcal{G}$ must have a $\big{(}\lfloor{\frac{\gamma+4}{2}}\rfloor\big{)}$-clique.

### v2

@@ 8.2 Consensus with an Edge Removal
These $\gamma$- and $(\gamma,\gamma)$ graphs achieve their respective maximum robustness levels using the fewest edges possible for a given number of nodes. In other words, removing any edge reduces their robustness levels at least by one. To validate this minimalistic property, we compare the consensus performance of the original $\gamma$- and $(\gamma,\gamma)$-MERGs with that of modified versions in which a single edge has been removed. For each graph, comparison is done using identical initial states for the normal agents and in the presence of $F$ misbehaving agents.
@@ 8.2.1 Simulations with $\gamma$-MERGs
We construct $5$-MERGs with $n=9$ and $n=10$ nodes as shown in Figure 2. Note these graphs achieve $5$-robustness with the fewest edges possible for a given number of nodes, and they tolerate at most $F=2$-local Byzantine agents. We set agents $i\in\{0,1\}$ as Byzantine, and agents $i\in\{2,\dots,n-1\}$ as normal. Each normal agent $i$ is initialized with a state $x_{i}[0]\in\mathbb{Z}_{\geq 0}$ sampled uniformly at random from the interval $[0,49]$ for $i\in\{2,3,4\}$ and from $[50,100]$ for $i\in\{5,\dots,n-1\}$. For all $t\geq 0$, all Byzantine agents adversarially share a value of $100$ with agents $i=0,\dots,\lceil\frac{n}{2}\rceil$ and of 100-100 with other remaining agents.
Figure 6 (a)-(b) and (e)-(f) illustrate the consensus results without and with edge removal. We have removed the edge $(3,8)$ for $n=9$ and $(4,9)$ for $n=10$ from the $5$-MERGs shown in Figure 2. In the $5$-MERGs, normal agents successfully reach resilient consensus (shown in Figure 6 (a) and (b)). However, removing the edge from $5$-MERG reduces its robustness to $r=4$, which is insufficient to guarantee resilient consensus in the presence of $F=2$ Byzantine agents. Thus, normal agents in the new graphs fail to reach resilient consensus (as shown by the red arrows in Figure 6 (e) and (f)).
@@ 8.2.2 Simulations with $(\gamma,\gamma)$-MERGs
We use Theorem 2 to construct $(5,5)$-MERGs with $n=9$ and $n=10$ agents, as shown in Figure 4. These graphs tolerate up to $F=4$-total malicious agents. We set agents $i\in\{0,\dots,3\}$ as malicious, and agents $i\in\{4,\dots,n-1\}$ as normal. Each normal agent $i$ is initialized with a state $x_{i}[0]\in\mathbb{R}$ sampled uniformly at random from the interval from $[0,49]$ for agents $i\in\{4,\dots,7\}$ and from $[50,100]$ for the other normal agents. Malicious agents share adversarial values: agent $3$ transmits 0 to its neighbors at all times, while the remaining malicious agents transmit $100$ to their neighbors at all times.
Figure 6 (c)-(d) and (g)-(h) illustrate the results without and with edge removal. We have removed the edge $(7,8)$ for $n=9$ and $(5,8)$ for $n=10$ from the $(5,5)$-MERGs visualized in Figure 4. In these $(5,5)$-MERGs, normal agents successfully reach resilient consensus (shown in Figure 6 (c) and (d)). However, removing the edge from each $(5,5)$-MERG reduces its robustness to at most $(5,4)$-robustness, which is insufficient to guarantee resilient consensus in the presence of $F=4$ malicious agents. As a result, normal agents fail to reach resilient consensus in the modified graphs (as shown by the red arrows in Figure 6 (g) and (h)).
@@ 10 Appendix
Here, the proof for Lemma 3 is provided below, accompanied by Figure 7 for illustrative support.
Figure 7: These figures demonstrate the first two iterations of the first part in the proof of Lemma 3. Throughout figures (a) to (d), we maintain $|\mathcal{S}_{1}|=\gamma-1$ and $|\mathcal{S}_{2}|=\gamma+1$. Figures (a) and (b) illustrate the first iteration. Here, $\mathcal{S}_{2}$ contains $\gamma+1$ nodes including $k$-clique $\mathcal{C}_{k,1}=\{v_{1},\cdots,v_{k}\}$ (colored in green), while $\mathcal{S}_{1}$ contains the remaining $\gamma-1$ nodes. Since $\mathcal{S}_{1}$ has to be $\gamma$-reachable, there exists a node $u_{1}\in\mathcal{S}_{1}$ that has edges with at least $\gamma$ nodes in $\mathcal{S}_{2}$. Figure (a) shows the case scenario where the yellow node $u_{1}$ does not have an edge with $v_{1}$ (highlighted in yellow). Then, $u_{1}$ forms a $k$-clique $\mathcal{C}_{k,2}$ with $\mathcal{C}_{k,1}\setminus\{v_{1}\}$. Consequently, in any future iteration, if any node $u_{i}\in\mathcal{S}_{1}$, $i\neq 1$, connects to at least $\gamma$ nodes in $\mathcal{S}_{2}$ but not to $v_{1}$, it will form a $(k+1)$-clique with $\mathcal{C}_{k,2}$. Then, $u_{1}$ gets swapped with a node in $S_{2}\setminus\mathcal{C}_{k,1}$, as shown in Figure (b). Figures (c) and (d) present the second iteration. Figure (c) presents the case scenario where $u_{2}$ only has $\gamma$ edges and does not have an edge with $v_{2}$ (highlighted in red). As before, if any node $u_{i}\in\mathcal{S}_{2}$, $i\neq 2$, connects to $\gamma$ nodes while missing an edge to $v_{2}$ in future iteration, it will form a $(k+1)$-clique. Then, $u_{2}$ gets swapped with a node in $\mathcal{S}_{2}\setminus U^{\prime}$ where $U^{\prime}=\mathcal{C}_{k,1}\cup\mathcal{C}_{k,2}$, as shown on Figure (d). The process continues, maintaining $\gamma$-reachability of $\mathcal{S}_{1}$, until a $\big(\lfloor\frac{\gamma+3}{2}\rfloor\big)$-clique is formed.
Proof.
@@
Second Part: Continuing from the previous paragraph, at $k=\lfloor\frac{\gamma+1}{2}\rfloor$ and for even values of $\gamma$, at worst the case 2) mentioned above is repeated $\frac{\gamma}{2}$ times. Then, $\mathcal{S}_{2}$ contains $\mathcal{K}=\{u_{1},\cdots,u_{\frac{\gamma}{2}}\}$, $\mathcal{C}_{k,1}=\{v_{1},\cdots,v_{\frac{\gamma}{2}}$}, and a node $p$ not in $\mathcal{K}$ and $\mathcal{C}_{k,1}$. Note both $\mathcal{K}$ and $\mathcal{C}_{k,1}$ disjointly form a $\big(\frac{\gamma}{2}\big)$-clique. We know that a node $u_{m}\in\mathcal{S}_{1}$, $m=\frac{\gamma+2}{2}$, must form a $m$-clique $\mathcal{Q}$ with either $\mathcal{K}$ or $\mathcal{C}_{k,1}$. We swap $u_{m}\in\mathcal{S}_{1}$ with $p\in\mathcal{S}_{2}$. Now let $\mathcal{P}=\mathcal{S}_{2}\setminus\mathcal{Q}$. Note that nodes in $\mathcal{P}$ form a $(\frac{\gamma}{2})$-clique such that $\mathcal{P}\cap\mathcal{Q}=\emptyset$. Since $\mathcal{S}_{1}$ is $\gamma$-reachable even after $u_{m}$ is swapped out, $\exists u_{m+1}\in\mathcal{S}_{1}$ such that $|\mathcal{N}_{u_{m+1}}\cap\mathcal{S}_{2}|\geq\gamma$. There are two cases:
@@
If 1), $\mathcal{Q}\cup\{u_{m+1}\}$ forms a $\big(\frac{\gamma+4}{2}\big)$-clique. If 2), $\mathcal{P}\cup\{u_{m+1}\}$ forms a $\big(\frac{\gamma+2}{2}\big)$-clique $\mathcal{P}^{\prime}$ such that $\mathcal{P}^{\prime}\cap\mathcal{Q}=\emptyset$. In this case, WLOG, let $\mathcal{P}^{\prime}\subset\mathcal{S}_{1}$ and $\mathcal{Q}\subset\mathcal{S}_{2}$ such that $|\mathcal{S}_{1}|=|\mathcal{S}_{2}|=\gamma$. Since either $\mathcal{S}_{1}$ or $\mathcal{S}_{2}$ is $\gamma$-reachable, all of the nodes in either $\mathcal{P}^{\prime}$ or $\mathcal{Q}$ have edges with one additional node in $\mathcal{S}_{2}$ or $\mathcal{S}_{1}$ respectively, forming a $\big(\frac{\gamma+4}{2}\big)$-clique. Thus, $\mathcal{G}$ must contain a $\big(\frac{\gamma+4}{2}\big)$-clique for even $\gamma$. Since $\mathcal{G}$ must have a clique size of $\frac{\gamma+3}{2}$ and $\frac{\gamma+4}{2}$ for odd and even $\gamma$ respectively, $\mathcal{G}$ must have a $\big(\lfloor{\frac{\gamma+4}{2}}\rfloor\big)$-clique. ∎